These are all separate questions. Thank you in advance.


A 500 mL base has a hydroxide concentration of 0.85 mol/L. If this base is diluted to 950 mL, what is the new hydroxide concentration of the diluted base?



In the science storage room, there is a stock solution for an unknown acid. This acid has a hydronium concentration of 0.55 mol/L, and there is 500mL of this acid left. If you dilute this acid to a volume of 1.3L, what is the hydronium concentration of the new acid?




What is the pOH of a solution with 8.5 g of HCl dissolved in 1.00 L of water.




What mass of NaOH is dissolved in 350mL of water if the pOH is 3.82




1
Expert's answer
2021-05-08T23:35:42-0400

"C_1V_1=C_2V_2"


"0.85\u00d7 0.5= 0.95C_2"

"C_2=0.45 mol \/L"




"0.55\u00d70.5= 1.3 C_2"

"C_2=0.21mol\/L"



"Molarity = \frac{moles}{volume}"


Moles of HCl = "\frac{8.5}{36.5}= 0.23 moles"


"Molarity=\frac{0.23}{0.01}=23 M"


"pOH = 14 - pH"


"pH= -log 23= - 1.36"


"pOH = 14 --1.36= 15.36"